OpenStudy (anonymous):

In an RC circuit, how long would it take, in multiples of time constant τ, for the capacitor to be 98% charged?

OpenStudy (anonymous):

Is R and C given?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4τ 0.9τ 9τ 90τ

OpenStudy (anonymous):

time constant = RC Q=q(1-e^(-t/RC)) Q/q = 1-e^(-t/τ) 1-0.98=e^(-t/τ) ln(1-0.98)= -t/τ t=3.91τ=4τ

OpenStudy (anonymous):

really thanks :)

OpenStudy (anonymous):

I need another help in my last que

OpenStudy (anonymous):

open a new post :D

OpenStudy (perl):

here is physics book and solutions http://thepiratebay.se/torrent/4885005/University_Physics_12th_edition_with_solutions

OpenStudy (anonymous):

it's all physics ?

OpenStudy (anonymous):

Thanks Perl for this book :D

OpenStudy (perl):

yeah i think so

OpenStudy (perl):

there might also be a college physics book around, this has integrals i believe

OpenStudy (anonymous):

really thanks :D